OpenStudy (anonymous):

a boat traveled 210 miles downstream and back. the trip downstream took 10 hours. the trip back took 70 hours. what is the speed of the boat stil in water? what is the speed of the current

OpenStudy (anonymous):

put r = rate of boat and say x = rate of stream then use distance = rate times time

OpenStudy (anonymous):

\[210=(r+x)\times 10\] that is rate with current is r + x times time \[210=(r-x)\times 70\] rate against current is r - x times time \[21=r+x\] \[3=r-x\] add the two equations to solve for r \[2r=24\] \[r=12\]

OpenStudy (anonymous):

so rate of boat is 12 and rate of current is 9 (since 12-9=3)

OpenStudy (anonymous):

i hope this is right so lets check. with the current you are going 12+9=21 miles per hour, and to go 210 miles at 21 mph takes 10 hours

OpenStudy (anonymous):

against the current is 12-9=3 mph and to go 210 miles at 3 mph takes 70 hours. looks good to me